OpenStudy (anonymous):

bouncing my answers off of someone please so i can make sure that i have the right answers

OpenStudy (anonymous):

Where are the answers?

OpenStudy (anonymous):

A tie pin which sells for $200.00 has a markup rate of 30% on the selling price. Find the amount of the markup and the cost My answer is: Markup Amount = $60.00, Cost = $150.00

OpenStudy (amistre64):

if 2 people have the same wrong answer, does it make it right?

OpenStudy (anonymous):

no

OpenStudy (amistre64):

100%+30% = 130% 200(1.30)

OpenStudy (anonymous):

260

OpenStudy (amistre64):

200/100 = 1% 2 = 1%; we want 30 of those 60 = 30% so that parts good

OpenStudy (amistre64):

then yes, add the markup to the original; and youre good

OpenStudy (anonymous):

so what i have is right

OpenStudy (amistre64):

60 and 260 would be right that 150 ill assume is a typo :)

OpenStudy (anonymous):

When a single card is drawn from a shuffled deck of cards, find the odds against getting a 9.

OpenStudy (amistre64):

as a side note: anything that is marked UP will not be sold for LESS

OpenStudy (anonymous):

my answer to this is 51/1 cause there is only 52 cards in a deck

OpenStudy (amistre64):

1 in 12 maybe?

OpenStudy (amistre64):

there are 4 9s which means there is 1 nine in 13 odds would be 1 to 12 then

OpenStudy (amistre64):

odds arent probabilities tho; they are weights

OpenStudy (amistre64):

1 for 12 against

OpenStudy (anonymous):

i did not unstand this one How many different three letter permutations can be formed from the letters in the word clipboard

OpenStudy (amistre64):

the word has nothing repeated so we just do 8 Pick 3

OpenStudy (amistre64):

8*9*7 i believe

OpenStudy (amistre64):

lol, i typoed it and came out right

OpenStudy (amistre64):

9P3 = 9*8*7

OpenStudy (anonymous):

i dont understand the question or the answer.. 8 pick 3?

OpenStudy (anonymous):

which is 504

OpenStudy (amistre64):

504 sounds good

OpenStudy (anonymous):

A pair of shoes with an original price of $300.00 is on sale for $210.00. Find the percent of the markdown

OpenStudy (amistre64):

9 options for the 1st letter, 8 options for the 2nd, and 7 options left for the 3rd

OpenStudy (amistre64):

9*8*7 is notated as 9P3

OpenStudy (anonymous):

my answer for this is :30%

OpenStudy (amistre64):

300/100 = 3 3 is 1% 3&70 = 210 so we are marked down 30% off

OpenStudy (amistre64):

3*70 that is

OpenStudy (anonymous):

210

OpenStudy (amistre64):

right, so we are paying 70% and are saving 30%; since 70+30 = 100

OpenStudy (anonymous):

so it is 30%

OpenStudy (anonymous):

loving

OpenStudy (anonymous):

sunny